# SMC-X: A Distributed Scalable Monte Carlo Simulation Method for Chemically Complex Alloys

### v1

@@ Abstract
To predict the complex chemical evolution in multicomponent alloys, it is highly desirable to have accurate atomistic simulation methods capable of reaching sufficiently large spatial and temporal scales. In this work, we advance the recently proposed SMC-X method through distributed computation on either GPUs or CPUs, pushing both spatial and temporal scales of atomistic simulation of chemically complex alloys to previously inaccessible scales. This includes a 16-billion-atom HEA system extending to the micrometer regime in space, and a 1-billion-atom HEA evolved over more than three million Monte Carlo swap steps, approaching the minute regime in time. We show that such large-scale simulations are essential for bridging the gap between experimental observations and theoretical predictions of the nanoprecipitate sizes in HEAs, based on analysis using the Lifshitz–Slyozov–Wagner (LSW) theory for diffusion-controlled coarsening. This work demonstrates the great potential of SMC-X for simulation-driven exploration of the chemical complexity in high-entropy materials at large spatial and temporal scales.
@@ 1 Introduction
Materials are composed of atoms, and the types of atoms together with their spatial arrangements govern a wide spectrum of physical properties, including mechanical, electrical, magnetic, and thermal behavior. However, the vast chemical complexity present in many technologically important materials poses formidable challenges for existing computational approaches. This challenge is particularly pronounced in high-entropy alloys (HEAs) 1, which contain multiple principal elements and therefore exhibit substantially increased chemical complexity. Moreover, this complexity is further compounded by the interplay and competition between order and disorder, as well as enthalpic and entropic effects, across diverse spatial, temporal, and temperature scales, making predictive modeling of such systems exceptionally demanding. Capturing these complexities, including chemical composition 2, disorder 3, chemical short-range order (SRO) 4, 5, chemical medium-range order (MRO) 6, multi-component intermetallic nanoparticles (MCINP) 7, and nano lamellae8, 9, requires modeling and simulating atomic systems over large spatial and temporal scales, with accuracy and generalizability of first principles method 10, 11, 12.
@@
In recent years, machine learning potentials (MLPs) 13, 14, 15 trained on DFT datasets 16, 17 have emerged as a highly promising solution to the aforementioned bottlenecks 18, 19. By fitting interatomic potential energy surfaces with machine learning models, MLPs can retain near-DFT accuracy 20 while dramatically increasing simulation efficiency and extending accessible system sizes 21. With an appropriate interaction cut-off radius, many MLPs can reduce the computational complexity of traditional DFT calculations from cubic scaling, $O(N^{3})$, to linear scaling, $O(N)$ 15, with respect to the number of atoms $N$. Moreover, by bypassing the explicit solution of the electronic structure, the computational prefactor can be reduced by orders of magnitude 22, enabling simulations at spatial scales previously deemed intractable 23. The integration of MLPs with molecular dynamics (MD), a foundational atomistic simulation technique renowned for its excellent scaling, has revolutionized computational materials science 24, chemistry 25, and biology 26 by enabling accurate, large-scale simulations at significantly reduced cost.
@@
Although the MLP + MD approach successfully scales simulations to unprecedented sizes 27, 28, 29, its intrinsic limitations restrict accessible time scales to the nanosecond regime, which typically requires millions of sequential MD steps. Consequently, it is extremely challenging, if not impossible, for MD to capture many important slow-evolving phenomena, such as precipitate formation in alloys 7. For instance, a single swap step between two neighboring lattice sites may take millions or even billions of MD steps to simulate. For such a purpose, Monte Carlo (MC) simulations offer unique advantages due to their characteristic of directly sampling configurational space, rather than tracking real-time trajectories as in MD 30, 31. However, compared to MD, the integration of conventional MC with MLP faces difficulties when scaled up to large systems 22. This challenge stems from the intrinsically sequential update nature in mainstream Monte Carlo algorithms, such as the Metropolis method, where trial moves are proposed site by site, inherently limiting parallelization 32. This bottleneck is in stark contrast with MD, where all atoms can be updated simultaneously at each step.
@@
Previous efforts to scale Monte Carlo (MC) simulations have introduced parallel trial moves using generalized checkerboard algorithms, with most implementations customized for specific interaction types or hardwares 33, 34, 35, 36, 37, 32, and only a few designed for general interactions 38, 22. Among all these implementations, the recently proposed scalable Monte Carlo at eXtreme (SMC-X) 22 method stands out due to its superb efficiency. Such a high efficiency is achieved via exposing the previously hidden degree of parallelism by dynamical link-cell (LC) decomposition and local interaction zone (LIZ), and by carefully mapping these degrees of parallelisms onto modern GPU architectures to remove unnecessary data movement. These innovations allow SMC-X to simulate systems as large as one billion atoms using a single GPU 22. This performance not only substantially exceeds all prior MC + MLP methods, but also exhibits advantages against state-of-the-art (SOTA) MD + MLP implementations in both the number of atoms and the atomistic simulation throughput, measured by $N_{\rm atom}\times N_{\rm steps}$ per second 22. A comparison of various computational methods for chemically complex alloys is shown in Tab. 1. It should be noted that while mesoscale methods such as the Potts model, kinetic Monte Carlo, and phase-fild methods can simulate very large spatial and temporal scales, they introduce coarse-grained approximations that rely on empirical parameters, making them less rigorous at the atomistic level, therefore are to compared in Tab. 1.
@@ 2 Theory and Methods
In this section, we present an introduction of the theoretical background of distributed SMC-X. Note that the basic concept of the SMC-X algorithm has been introduced in Ref. 22, but the focus was on its application in materials science. Here we present a more detailed discussion and introduce the distributed implementation on both CPUs and GPUs.
@@ 2.1 Monte Carlo simulation
$\langle\hat{O}\rangle=\sum_{k}\frac{O(\sigma_{k})e^{-\beta H(\bm{\sigma}_{k})}}{Z}.$ (2)
A sample following the probability distribution of Eq. 1 can be generated efficiently via Markov chain Monte Carlo (MCMC) simulation. For such a purpose, the Metropolis-Hastings algorithms is the most commonly used method. The Metropolis-Hasting algorithm starts with proposing a move (trial) at each lattice site, with the acceptance probability of this trial given by:
@@
The trial moves within the Metropolis-Hasting algorithm are inherently sequential, which poses a significant challenge to parallelization. Although advanced techniques such as cluster MC 39 or the Wang–Landau algorithms 40, 41, 42 can alleviate this limitation by reducing or avoiding sequential site-by-site updating, they are often restricted in scope and lack general applicability. A more general approach leverages the assumption of short-range interactions: If interactions are sufficiently localized, then spatially distant regions can be updated independently and in parallel. In the simple Ising model, this principle underlies the well-known checkerboard algorithm. Extensions of this approach to three-dimensional lattices and beyond nearest-neighbor interactions are conceptually straightforward for specific models 33. However, generalization to arbitrary short-range interactions is more limited. One notable example is the Scalable Parallel Monte Carlo (SPMC) algorithm 38, as implemented in the LAMMPS package 43. Despite its utility, SPMC has disadvantage for further scaling up: to minimize the weak spatial correlations in SPMC algorithm, the total volume of the domains need to be maximized 38, which underutilizes the available parallelism, especially on architectures with a high degree of concurrency. As a result, current application of SPMC are limited systems of less than a million atoms 29, 44.
Figure 2: A schematic to illustrate the SMC-X algorithm, including link-cell (LC), local interaction zone (LIZ), and domain decomposition scheme, using the 2D square lattice as an example. LIZ+ represents sites needed for calculating the energy change due to possible MC trial. Adapted from Ref.  22
@@ 2.2 Distributed SMC-X
In the SMC-X framework 22, the challenge of site-by-site sequential updating is addressed through two key innovations: dynamical link-cell (LC) and local interaction zone (LIZ), as illustrated in Fig. 2. Compared to SPMC 38, the LCs are dynamically constructed to ensure that every site is visited exactly once per MC step, and that the detailed balance condition is strictly satisfied. The resulting degree of parallelism is given by:
@@
where $N_{atom}$ is the total number of atoms in the supercell, and $N_{LC}$ is the number of atoms within LC. By choosing the smallest LC size that fully separates short-range interactions, the parallelism $P_{\mathrm{LC}}$ can be maximized. The LIZ is introduced to harness the short-range interaction to reduce the computational complexity of a MC trial move from $O(N)$ to $O(1)$. Furthermore, by computing the atomic features within each LIZ on the fly and tightly couple it with the MC simulation, the memory usage in SMC-X can be drastically reduced, and near-core caches can be effectively utilized to significantly enhance the computational throughput.
In addition to LC and LIZ, there are additional parallel opportunities in SMC-X, as illustrated in the algorithm part in Fig. 1. Among them, the lattice and the temperature parallelism can be implemented as MPI ranks for distributed computation over multiple chips. In lattice parallelism, the complete lattice is partitioned across multiple GPUs, with a ghost layer used to exchange boundary information between neighboring sublattices, as illustrated in Fig. 2. Under this scheme, if the number of atoms per GPU remains fixed, the total system size then increases proportionally to the number of GPUs, demonstrating weak scaling. Conversely, when the same system is distributed over more GPUs, the time required for a single MC step decreases, enabling more steps to be simulated within the same wall-clock time, an indication of strong scaling. Therefore, the strong and weak scaling behaviors in lattice parallelism represent the potential to push the atomistic simulation to larger spatial and temporal scales, respectively, by simply adding more computational resources.
The temperature parallelism can be discussed in two different situations. First, note that in the canonical ensemble, temperature is fixed, and simulations at different temperatures are essentially independent. Therefore, the MC simulation is embarrassingly parallel when properties at different temperatures are needed, such as to search for an order-disorder transition temperature. On the other hand, it is well known that the convergence speed of conventional MC can be slow, particularly near second-order transition, and the replica-exchange (RE, also known as parallel tempering) is a widely used technique to address this problem. In RE, multiple replicas of the same system are simulated concurrently at different temperatures. Configurations between replicas are periodically exchanged according to a Metropolis criterion that preserves detailed balance. The transition probability for two configurations with energies $E_{i}$ and $E_{j}$ is given by:
@@
$T_{i}$ and $T_{j}$ are the temperatures for configuration $i$ and $j$. By exchanging configurations between neighboring temperatures, the replica-exchange algorithm mitigates the risk of the system becoming trapped in local energy minima. This is especially critical near second-order phase transition points 39, 40, where the divergence of the correlation length makes it challenging for simulations at a fixed temperature to escape metastable states. For large systems, an additional complication needs to be considered: the energy fluctuation decreases as the system size increases. Consequently, the energy distributions of adjacent replicas may overlap insufficiently, leading to low exchange rates. To ensure reasonable exchange rate, the temperature difference $\Delta T$ need to be adjusted in accordance with the system size $N_{\text{atom}}$ as:
@@
For large atomic systems, the required temperature spacing $\Delta T$ becomes very small (e.g., 0.1 K for 1 billion atoms), which significantly increases the number of replicas and thus the overall computational cost. Limited by computing resources, we do not attempt a full RE simulation in this work. Note that the $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ HEA studied in this work does not show obvious order-disorder transition in its $C_{v}$ curve, as illustrated in Ref. 22.
@@ 3 Results and discussion
In this section, we present the simulation results for the $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ HEA. Two models, EPI and qSRO, are used. The EPI model31, 45 contains only pairwise interactions, which is computationally highly efficient since there is no need to explicitly evaluate the local energy changes within the LIZ. On the other hand, the qSRO model 22 represents a general nonlinear short-range model, where the local energies of all lattice sites within LIZ are calculated to determine the energy change due to a MC move. For SMC-GPU, the computational platform is a 2-node GPU cluster. Each computational node contains 8 80-GB NVIDIA H800 GPUs interconnected via 400 GB/s NVLink, and the 2 nodes are connected with a 200 GB/s InfiniBand network. For SMC-CPU, the platform is a node with 2 Kungpeng ARM CPUs. Each CPU contains multiple NUMA domains, with dozens of physical cores in each domain.
The DFT dataset is calculated with the LSMS method 46. LSMS is an all-electron electronic structure calculation method in which the computational cost scales linearly with respect to the number of atoms. For FeCoNiAlTi, we use a 100 atom supercell with a lattice constant of 6.72 Bohr (0.356 nm). We employ a spin-polarized scheme to account for the magnetic interactions in the system. The angular momentum cutoff $l_{max}$ for the electron wavefunctions is chosen as 3, and the LIZ is chosen as 86 atoms. We used PBE as the exchange-correlational functional. To enhance the representativeness of the DFT data, the total dataset is made up of 120 randomly generated configurations and 100 configurations from Monte Carlo simulation, as proposed in 31. Therefore, the total number of configurations is 220, with each comprising 100 atoms.

### v2

@@ Abstract
To predict the complex chemical evolution in multicomponent alloys, it is highly desirable to have accurate atomistic simulation methods capable of reaching sufficiently large spatial and temporal scales. In this work, we advance the recently proposed SMC-X method through distributed computation on either GPUs or CPUs, pushing both spatial and temporal scales of atomistic simulation of chemically complex alloys to previously inaccessible scales. This includes a record-breaking 128-billion-atom HEA system extending to the micrometer regime in space, and a 1-billion-atom HEA evolved over more than three million Monte Carlo swap steps, approaching the minute regime in time. We show that such large-scale simulations are essential for bridging the gap between experimental observations and theoretical predictions of the nanoprecipitate sizes in HEAs, based on analysis using the Lifshitz–Slyozov–Wagner (LSW) theory for diffusion-controlled coarsening. This work demonstrates the great potential of SMC-X for simulation-driven exploration of the chemical complexity in high-entropy materials at large spatial and temporal scales.
@@ 1 Introduction
Materials are composed of atoms, and the types of atoms together with their spatial arrangements govern a wide spectrum of physical properties, including mechanical, electrical, magnetic, and thermal behavior. However, the vast chemical complexity present in many technologically important materials poses formidable challenges for existing computational approaches. This challenge is particularly pronounced in high-entropy alloys (HEAs) EGeorge_Nature, which contain multiple principal elements and therefore exhibit substantially increased chemical complexity. Moreover, this complexity is further compounded by the interplay and competition between order and disorder, as well as enthalpic and entropic effects, across diverse spatial, temporal, and temperature scales, making predictive modeling of such systems exceptionally demanding. Capturing these complexities, including chemical composition ChemicalComposition2019Nature, disorder AtomComplexity, chemical short-range order (SRO) SRO_NCS_2023, APT_CoCrNi_NatureMat_2024, chemical medium-range order (MRO) mediumRangeOrderNC, multi-component intermetallic nanoparticles (MCINP) Yang933, and nano lamellaeCTLiuNC, MaxPlankNC2025, requires modeling and simulating atomic systems over large spatial and temporal scales, with accuracy and generalizability of first principles method LIU2023101018, LIU2019107955, Eisenbach_2019.
@@
In recent years, machine learning potentials (MLPs) JCTCBP, Zhang_2025, MPNNs_2023 trained on DFT datasets OQMD, Horton2025 have emerged as a highly promising solution to the aforementioned bottlenecks UniversalModel_NCS, 2025Nature. By fitting interatomic potential energy surfaces with machine learning models, MLPs can retain near-DFT accuracy Ong_JPC_2020 while dramatically increasing simulation efficiency and extending accessible system sizes doi:10.1126/science.abn3445. With an appropriate interaction cut-off radius, many MLPs can reduce the computational complexity of traditional DFT calculations from cubic scaling, $O(N^{3})$, to linear scaling, $O(N)$ MPNNs_2023, with respect to the number of atoms $N$. Moreover, by bypassing the explicit solution of the electronic structure, the computational prefactor can be reduced by orders of magnitude liuNPJ2025, enabling simulations at spatial scales previously deemed intractable 10.1145/3458817.3487400. The integration of MLPs with molecular dynamics (MD), a foundational atomistic simulation technique renowned for its excellent scaling, has revolutionized computational materials science 2025Nature, chemistry Zhang2024, and biology Gkeka2020 by enabling accurate, large-scale simulations at significantly reduced cost.
@@
Although the MLP + MD approach successfully scales simulations to unprecedented sizes 10.1145/3581784.3627041, 2020NPJ_Li, Yin2021, its intrinsic limitations restrict accessible time scales to the nanosecond regime, which typically requires millions of sequential MD steps. Consequently, it is extremely challenging, if not impossible, for MD to capture many important slow-evolving phenomena, such as precipitate formation in alloys Yang933. For instance, a single swap step between two neighboring lattice sites may take millions or even billions of MD steps to simulate. For such a purpose, Monte Carlo (MC) simulations offer unique advantages due to their characteristic of directly sampling configurational space, rather than tracking real-time trajectories as in MD Korman_npj, LIU2021110135. However, compared to MD, the integration of conventional MC with MLP faces difficulties when scaled up to large systems liuNPJ2025. This challenge stems from the intrinsically sequential update nature in mainstream Monte Carlo algorithms, such as the Metropolis method, where trial moves are proposed site by site, inherently limiting parallelization Anderson2016SMC. This bottleneck is in stark contrast to that in MD, where all atoms can be updated simultaneously at each step.
@@
Previous efforts to scale Monte Carlo (MC) simulations have introduced parallel trial moves using generalized checkerboard algorithms, with most implementations customized for specific interaction types or hardware PREIS20094468, BLOCK20101549, ROMERO2020107473, liu2019high, OrtegaZamorano2013FPGAIsing, Anderson2016SMC, and only a few designed for general interactions sadigh2012scalable, liuNPJ2025. Among all these implementations, the recently proposed scalable Monte Carlo at eXtreme (SMC-X) liuNPJ2025 method stands out due to its superb efficiency. Such a high efficiency is achieved via exposing the previously hidden degree of parallelism by dynamical link-cell (LC) decomposition and local interaction zone (LIZ), and by carefully mapping these degrees of parallelisms onto modern GPU architectures to remove unnecessary data movement. These innovations allow SMC-X to simulate systems as large as one billion atoms using a single GPU liuNPJ2025. This performance not only substantially exceeds all prior MC + MLP methods, but also exhibits advantages against state-of-the-art (SOTA) MD + MLP implementations in both the number of atoms and the atomistic simulation throughput, measured by $N_{\rm atom}\times N_{\rm steps}$ per second liuNPJ2025. A comparison of various computational methods for chemically complex alloys is shown in Tab. 1. It should be noted that while mesoscale methods such as the Potts model, kinetic Monte Carlo, and phase-fild methods can simulate very large spatial and temporal scales, they introduce coarse-grained approximations that rely on empirical parameters, making them less rigorous at the atomistic level, therefore are not listed in Tab. 1.
@@ 2 Theory and Methods
In this section, we present an introduction of the theoretical background of distributed SMC-X. Note that the basic concept of the SMC-X algorithm has been introduced in Ref. liuNPJ2025, but the focus was on its application in materials science. Here we present a more detailed discussion and introduce the distributed implementation on both CPUs and GPUs.
@@ 2.1 Monte Carlo simulation
$\langle\hat{O}\rangle=\sum_{k}\frac{O(\sigma_{k})e^{-\beta H(\bm{\sigma}_{k})}}{e^{-\beta H(\bm{\sigma}_{k})}}.$ (2)
A sample following the probability distribution of Eq. 1 can be efficiently generated via a Markov chain Monte Carlo (MCMC) simulation. For such a purpose, the Metropolis-Hastings algorithms is the most commonly used method. The Metropolis-Hasting algorithm starts with proposing a move (trial) at each lattice site, with the acceptance probability of this trial given by:
@@
The trial moves within the Metropolis-Hasting algorithm are inherently sequential, which poses a significant challenge to parallelization. Although advanced techniques such as cluster MC PhysRevLett.58.86 or the Wang–Landau algorithms PhysRevLett.86.2050, 10.1145/1654059.1654125, PhysRevE.97.043301 can alleviate this limitation by reducing or avoiding sequential site-by-site updating, they are often restricted in scope and lack general applicability. A more general approach leverages the assumption of short-range interactions: If interactions are sufficiently localized, then spatially distant regions can be updated independently and in parallel. In the simple Ising model, this principle underlies the well-known checkerboard algorithm. Extensions of this approach to three-dimensional lattices and beyond nearest-neighbor interactions are conceptually straightforward for specific models PREIS20094468. However, generalization to arbitrary short-range interactions is more limited. One notable example is the Scalable Parallel Monte Carlo (SPMC) algorithm sadigh2012scalable, as implemented in the LAMMPS package THOMPSON2022108171. Despite its utility, SPMC has disadvantage for further scaling-up: to minimize the weak spatial correlations in SPMC algorithm, the total volume of the domains need to be maximized sadigh2012scalable, which underutilizes the available parallelism, especially on architectures with a high degree of concurrency. As a result, current applications of SPMC are limited systems of less than a million atoms Yin2021, YAO2024120457.
Figure 2: A schematic to illustrate the SMC-X algorithm, including link-cell (LC), local interaction zone (LIZ), and domain decomposition scheme, using the 2D square lattice as an example. LIZ+ represents sites needed for calculating the energy change due to possible MC trial. Adapted from Ref.  liuNPJ2025
@@ 2.2 Distributed SMC-X
In the SMC-X framework liuNPJ2025, the challenge of site-by-site sequential updating is addressed through two key innovations: dynamical link-cell (LC) and local interaction zone (LIZ), as illustrated in Fig. 2. Compared to SPMC sadigh2012scalable, the LCs are dynamically constructed to ensure that every site is visited exactly once per MC step, and that the detailed balance condition is strictly satisfied. The resulting degree of parallelism is given by:
@@
where $N_{atom}$ is the total number of atoms in the supercell, and $N_{LC}$ is the number of atoms within LC. By choosing the smallest LC size that fully separates short-range interactions, the parallelism $P_{\mathrm{LC}}$ can be maximized. The LIZ is introduced to exploit the short-range interaction to reduce the computational complexity of an MC trial move from $O(N)$ to $O(1)$. Furthermore, by computing the atomic features within each LIZ on the fly and tightly coupling it with the MC simulation, the memory usage in SMC-X can be drastically reduced, and near-core caches can be effectively utilized to significantly enhance computational throughput.
In addition to LC and LIZ, there are additional parallel opportunities in SMC-X, as illustrated in the algorithm part in Fig. 1. Among them, the lattice and the temperature parallelism can be implemented as MPI ranks for distributed computation over multiple chips. In lattice parallelism, the complete lattice is partitioned across multiple GPUs, with a ghost layer used to exchange boundary information between neighboring sublattices, as illustrated in Fig. 2. Under this scheme, if the number of atoms per GPU remains fixed, the total system size increases proportionally to the number of GPUs, demonstrating weak scaling. Conversely, when the same system is distributed over more GPUs, the time required for a single MC step decreases, enabling more steps to be simulated within the same wall-clock time, an indication of strong scaling. Therefore, the strong and weak scaling behaviors in lattice parallelism represent the potential to push the atomistic simulation to larger spatial and temporal scales, respectively, by simply adding more computational resources.
The temperature parallelism can be discussed in two different situations. First, note that in the canonical ensemble, temperature is fixed, and simulations at different temperatures are essentially independent. Therefore, the MC simulation is embarrassingly parallel when properties at different temperatures are needed, such as to search for an order-disorder transition temperature. On the other hand, it is well known that the convergence speed of conventional MC can be slow, particularly near second-order transition, and replica-exchange (RE, also known as parallel tempering) is a widely used technique to address this problem. In RE, multiple replicas of the same system are simulated concurrently at different temperatures. Configurations between replicas are periodically exchanged according to a Metropolis criterion that preserves detailed balance. The transition probability for two configurations with energies $E_{i}$ and $E_{j}$ is given by:
@@
$T_{i}$ and $T_{j}$ are the temperatures for configuration $i$ and $j$. By exchanging configurations between neighboring temperatures, the replica-exchange algorithm mitigates the risk of the system becoming trapped in local energy minima. This is especially critical near second-order phase transition points PhysRevLett.58.86, PhysRevLett.86.2050, where the divergence of the correlation length makes it challenging for simulations at a fixed temperature to escape metastable states. For large systems, an additional complication needs to be considered: the energy fluctuation decreases as the system size increases. Consequently, the energy distributions of adjacent replicas may overlap insufficiently, leading to low exchange rates. To ensure reasonable exchange rate, the temperature difference $\Delta T$ need to be adjusted in accordance with the system size $N_{\text{atom}}$ as:
@@
For large atomic systems, the required temperature spacing $\Delta T$ becomes very small (e.g., 0.1 K for 1 billion atoms), which significantly increases the number of replicas and thus the overall computational cost. Limited by computing resources, we do not attempt a full RE simulation in this work. Note that the $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ HEA studied in this work does not show obvious order-disorder transition in its $C_{v}$ curve, as illustrated in Ref. liuNPJ2025.
@@ 3 Results and discussion
In this section, we present the simulation results for the $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ HEA. Two models, EPI and qSRO, are used. The EPI modelLIU2021110135, ZHANG2020108247 contains only pairwise interactions, which is computationally highly efficient since there is no need to explicitly evaluate the local energy changes within the LIZ. On the other hand, the qSRO model liuNPJ2025 represents a general nonlinear short-range model, where the local energies of all the lattice sites within LIZ are calculated to determine the energy change due to a MC move. For SMC-GPU, the computational platform is a 2-node GPU cluster. Each computational node contains 8 80-GB NVIDIA H800 GPUs interconnected via 400 GB/s NVLink, and the 2 nodes are connected with a 200 GB/s InfiniBand network. For SMC-CPU, the platform is a node with 2 Kungpeng ARM CPUs. Each CPU contains multiple NUMA domains, with dozens of physical cores in each domain.
The DFT data-set is calculated with the LSMS method PhysRevLett.75.2867. LSMS is an all-electron electronic structure calculation method in which the computational cost scales linearly with respect to the number of atoms. For FeCoNiAlTi, we use a 100 atom supercell with a lattice constant of 6.72 Bohr (0.356 nm). We employ a spin-polarized scheme to account for the magnetic interactions in the system. The angular momentum cutoff $l_{max}$ for the electron wavefunctions is chosen as 3, and the LIZ is chosen as 86 atoms. We used PBE as the exchange-correlational functional. To enhance the representativeness of the DFT data, the total dataset is made up of 120 randomly generated configurations and 100 configurations from Monte Carlo simulation, as proposed in LIU2021110135. Therefore, the total number of configurations is 220, with each comprising 100 atoms.